# Spatialyze: A Geospatial Video Analytics System with Spatial-Aware Optimizations

### v3

@@ 5.2.2. Video Processor
[…]
Figure 4. Baseline Video Processing Plan
Executing the plan computes 3D tracks of objects in a video. However, streaming each video through all operators is computationally expensive and not always necessary. Based on the users’ filter predicates from when they Filter the World, Spatialyze only includes the necessary operators in the execution plan. For instance, a predicate based on object types like obj.type=="car" requires objects’ types and hence the execution plan will include lines LABEL:line:plan-decode and LABEL:line:plan-detect. Likewise, a predicate with distance or contain functions requires objects’ 3D location and therefore include lines LABEL:line:plan-decode, LABEL:line:plan-detect, and LABEL:line:plan-3d in the plan. Finally, a predicate with a headingDiff function requires objects’ moving directions and include all lines in the plan.
We implemented each optimization technique in § 6 as a streaming operator. Two of the optimization operators execute users’ predicate early on to reduce the number of input (video frames and objects) into later operators. We designed our video processor that constructs plans using streaming operators; therefore, future optimization techniques and video processing functions can be added, addressing the runtime efficiency and extensibility challenges.
5.2.3. Movable Objects Query Engine
This stage takes in users’ filter predicates from their workflow’s Filter and resulting Movable Objects from the video processor. It then filters the Movable Objects with the users’ predicates in 2 steps. First, it streams the Movable Objects into a newly created table in the geospatial metadata store. The table has a spatial index on objects’ tracks and a temporal index on the time period that the objects exist. Then, it translates the user’s filter predicates into an SQL query and execute it in the geospatial metadata store against the Movable Objects, Cameras, and RoadNetwork. When the user’s predicates involve multiple objects (e.g., distance(car1,car2)<5), Spatialyze self-joins the Movable Objects table to find all objects that coexist in a given time period using the temporal index to speed up the join. When the predicates involve spatial relationships between a Movable Object and a RoadNetwork (e.g., contains(road,car)), Spatialyze uses the the spatial index to speed up the joins between the Movable Objects and RoadNetwork tables. We implemented the metadata store using MobilityDB (Zimányi et al., 2020) as it provides necessary spatial and temporal index for movement objects. The index along with its spatiotemporal data type are suited for representing and processing queries on Movable Objects and Geographic Constructs.
5.2.4. Output Composer
This stage composes the Movable Objects query results to a users’ preferred format for them to observe. Calling getObjects, Spatialyze returns a list of Movable Objects. Calling saveVideos, Spatialyze encodes the frames containing the Movable Objects into video files with colored bounding boxes annotation.
Fig. 2 shows the execution stages for the workflow shown in Fig. 1. It contains all stages mentioned above. A user first build the world by adding data. Spatialyze’s Data Integrator creates an indexed table consisting of RoadNetwork and joins the Video and its corresponding Camera. Then, Spatialyze’s Video Processor creates an optimized video processing plan with 4 processing operators and 4 optimization operators (to be discussed in § 6). Parts of the users’ predicates are executed in Road Visibility Pruner and Object Type Pruner. The Movable Objects Query Engine streams the Movable Objects from Video Processor into the geospatial metadata store and filters them with the users’ predicates. Finally, as the workflow asks for videos, Spatialyze formats and saves the frames that have the filtered objects into video snippet files.
6. Video Processing Optimization
SOTA video inference uses ML models to extract the desired information, such as object positions and their trajectories within each video. Running such models is computationally expensive. Spatialyze’s video processor comes with 4 optimization techniques to reduce such runtime bottleneck. To be discussed below, the idea behind our optimization is to leverage the geospatial properties of the videos, and the Movable Objectsinherited physical behaviors. We implement each optimization as a streaming operator in the Video Processing stage in § 5.2.2. Besides the existing 4 processing operators, Spatialyze may place optimization operators in the middle of the video processing plan to reduce the amount of input into later processing operators based on the filter predicates. In § 6.1, Spatialyze places the Road Visibility Pruner after the Video Decoder, when users specify the Geographic Constructs the objects of interest must be in. Doing so reduces the number of input frames for the rest of the operators in the plan. In § 6.2, If users filter on object types, then Spatialyze places the Object Type Pruner right after the Object Detector. It reduces the number of objects that Spatialyze needs to calculate for their 3D locations and tracking. In § 6.3, Spatialyze replaces the ML-based 3D Location Estimation operator with the much faster 0pt, if Spatialyze can assume that the objects of interest are on the ground, such as cars or people. In § 6.4, Spatialyze places the Exit Frame Sampler in between the 3D Location Estimator and the Object Tracker, if the users’ predicates focus on cars. It utilizes the 3D locations of the objects and the user-provided metadata to reduce the number of frames that the Object Tracker processes.
Fig. 3 shows where the video processor executes all the optimization steps. Line LABEL:line:pred-contain in Fig. 1 corresponds to Road Visibility Pruner. Line LABEL:line:pred-type filters by object type, so Spatialyze adds Object Type Pruner. The objects of interest are cars or trucks, so Spatialyze adds Exit Frame Sampler and 0pt. In the following sections, we discuss each optimization in detail.
6.1. Road Visibility Pruner
In our experiments, the runtime of ML models (Jocher et al., 2022; Zhou et al., 2019; Du et al., 2023) in the video processing plan takes 90% of the entire workflow execution runtime, on average. Hence, the goal of the Road Visibility Pruner is to remove frames that do not contain objects of user’s interest to reduce the overall runtime of the video processing plan.
6.1.1. High-Level Concepts
The Road Visibility Pruner uses Geographic Construct’s visibility as a proxy for a Movable Object’s visibility. Specifically, a predicate includes contains(road,obj) and distance(cam,obj)<d means that obj is not visible in the frames where road is not visible within d meters; therefore, Spatialyze prunes out those frames. In Fig. 1, the Road Visibility Pruner partially executes lines LABEL:line:pred-contain where the user searches for “cars within 50 meters at an intersection.” Spatialyze removes all frames that do not contain visible intersection within 50 meters as these frames will not contain any car of users’ interest.
The Road Visibility Pruner only concerns the visibility of Geographic Constructs. Therefore, it uses Geographic Constructs and Camera to decide whether to prune out a video frame. It uses the intrinsic and extrinsic properties of Camera (Zhang, 2014) to determine the viewable area of each of the camera according to the world coordinate system. [^2] World Coordinate System is a 3 dimensional Euclidean space, representing a World. It then determines if the viewable area of each frame includes any Geographic Construct of interest; frames that do not contain such constructs are dropped and will not be processed in the later processing operators. Since the Road Visibility Pruner will be the first processing operator if applied, it can reduce the number of frames that need to be processed by subsequent ML models in the plan, and hence can substantially reduce overall execution time.
6.1.2. Algorithm
The Road Visibility Pruner works on a per-frame basis. For each video frame $i$ shot by a Camera at location $l_{i}$ with rotation $r_{i}$ and camera intrinsic $it_{i}$, the Road Visibility Pruner performs three steps to determine if the frame should be pruned out. First, it computes the 3D viewable space of the Camera at frame $i$. Then, it uses the viewable space to determine Geographic Constructs that are visible to the Camera at frame $i$. Finally, it uses the users’ filter predicates to determine if there is any Geographic Construct of interest visible in the frame. We explain each step below.
First, the Road Visibility Pruner calculates the 3D viewable space of the Camera at frame $i$. The viewable space is a pyramid (Fig. 5 side view). A⃝ is the location of the camera. B⃝, C⃝, D⃝, and E⃝ represent the top-left, top-right, bottom-right, and bottom-left of the frame at $d$ meters in front of the camera. We compute B⃝, C⃝, D⃝, and E⃝ by converting the mentioned 4 corner points of the frame from pixel (2D) to world (3D) coordinates. This is done in 3 steps by converting a point from: 1) Pixel to Camera Coordinate System; 2) Camera to World Coordinate System; and 3) Pixel to World Coordinate System.
@@ 6.1.2. Algorithm
Pixel to Camera Coordinate System. The equation for converting a 3D point $(x_{c},y_{c},z_{c})$ from camera coordinate system333 Camera Coordinate System is a 3 dimensional Euclidean space from the perspective of a camera. The origin point is at the camera position. Z-axis and X-axis point toward the front and right of the camera. Y-axis points downward from the camera. to a 2D point $(x_{p},y_{p})$ in pixel coordinate system is:
@@
where $it_{i}$ represents the camera intrinsic, a set of parameters representing Camera’s focal lengths ($f_{x},f_{y}$), skew coefficient ($s$), and optical center in pixels ($x_{0},y_{0}$). The intrinsic is specific to each camera, independent of its location and rotation. It is used to convert 3D points from camera perspective to 2D points in pixels.
@@
Camera to World Coordinate System. For this conversion, we set up $R$, a $3\times 3$ matrix, representing the 3D rotation of the camera that is derived from the camera’s rotation quaternion. And, $t$ is a 3D vector that represents the camera’s translation. Then, a $3\times 4$ extrinsic matrix $[R|t]$ converts 3D points from camera coordinate system $(x_{c},y_{c},z_{c})$ to 3D points $(x,y,z)$ in world coordinate system:
Pixel to World Coordinate System. Combining the previous 2 conversions in Eq. 3 and Eq. 4, we can convert 2D points from pixel coordinate system to 3D points in world coordinate system using:
4 corners of a video frame in World Coordinate System. To convert the 4 corners of a video frame from pixel to world coordinates, we modify Eq. 5 where we substitute $z_{c}$ with $d$, and $x_{p}$ and $y_{p}$ with actual pixel values representing the 4 corners of the video frame; $w$ and $h$ are the width and height in pixels of the video frame, and each of $B,C,D,E$ is a $3\times 1$ vector representing its 3D location.
After the corners of each frame are translated to world coordinates, the Road Visibility Pruner uses the viewable space to determine the Geographic Constructs that are visible by Camera at frame $i$. Each Geographic Construct’s area is defined using a 2D polygon. Specifically, all polygons representing Geographic Constructs are 2D in the $z=0$ plane. Therefore, the Road Visibility Pruner projects the computed 3D viewable space onto the $z=0$ plane. We define the top-down 2D viewable area as the convex hull (Sklansky, 1982) of the projected vertices A⃝, B⃝, C⃝, D⃝, and E⃝ (yellow highlight in Fig. 5’s top view). We then spatially join the 2D viewable area with Geometric Constructs in the geospatial metadata store, making use of their geospatial indexing created in § 5.2.1. Any Geometric Construct that overlaps with the 2D viewable area is considered visible to the Camera at frame $i$. Once computed, we get visibleGTypes, a set of Geometric Construct types that are visible to the Camera at frame $i$.
Finally, the Road Visibility Pruner uses the users’ filter predicates to determine if there is any Geographic Construct of interest visible in the frame. Examining the contains predicates in the filter, it assigns each contains to $True$ if its first argument is in visibleGTypes and $False$ otherwise. If the value of the transformed filter predicates is $False$, then the Road Visibility Pruner prunes out that frame.
Figure 5. (Side) A 3D viewable space of a camera in pyramid shaped. (Top) Projected area of the pyramid onto $z=0$ plane.
@@ 6.1.3. Benefits and Limitations
The Road Visibility Pruner has an insignificant overhead runtime compared to the runtime it can save. Our experiments show that the Road Visibility Pruner takes 0.1% of the video processing runtime to execute, while saving up to 19.6% of the video processing runtime. So in the worst case that Road Visibility Pruner cannot prune out any video frames, the runtime increase is still negligible. However, having the Road Visibility Pruner in the video processing plan may cause tracking accuracy to drop in some scenarios. Using § 3 as an example, if a car starts at an intersection, exits the intersection, and enters an intersection again. During the time that the car is not in any intersection, the camera may not capture any frames containing intersection. The Road Visibility Pruner will prune out the period of video frames where there is no intersection visible to the camera. This can lead to the object tracker considering the car before leaving the intersection and the car after entering the intersection for the second time to be two different cars.
Figure 6. In Fig. 1, the video processor executes expensive ML models only on video frames with a visible intersection.
@@ 6.2.1. High-Level Concepts
Spatialyze only needs to track objects that are of interest as stated in the workflow. The emphObject Type Pruner prunes out irrelevant objects from being tracked. We analyze users’ filter predicates to look for types of objects that are necessary for computing the workflow outputs. Using the predicates in Fig. 1 as an example, the Object Type Pruner partially executes the line LABEL:line:pred-type, where users only look for cars and trucks in the workflow outputs. The trajectories of other objects (e.g., humans) will not appear in the final workflow outputs anyway, so tracking them is unnecessary. Therefore, we only need to input objects that are detected as cars or trucks into object tracker to reduce its workload.
@@ 6.2.2. Benefits and Limitations
The Object Type Pruner primarily reduces object tracker’s runtime. Nevertheless, the Object Type Pruner has low overhead runtime of 0.06% and does not require any geospatial metadata, but can reduce up to 69% of the runtime spent on object tracking, which is 26% of the video processor runtime.
@@
Spatialyze uses an object’s distance from camera and its calculated bounding box to estimate its 3D location in our video processing plan. Specifically, we use YOLOv5 (Jocher et al., 2022) to estimate the bounding box and Monodepth2 (Godard et al., 2019) to estimate the object’s distance from camera in our baseline plan. Monodepth2 only requires monocular images (videos) to estimate each object’s distance from camera; therefore, it is extremely flexible in general use cases. Spatialyze can always use Monodepth2 to estimate any object’s distance from camera in a video frame. However, this algorithm is slow and can take upto 48% portion of the total baseline video processing runtime. Therefore, we explored an alternative algorithm that leverage existing geospatial metadata to speed up the process for estimating 3D object location.
@@ 6.3.1. High-Level Concepts
0pt leverages Camera to estimate an object’s 3D locations from its 2D bounding box. It is a substitution for Monodepth2 to estimate objects’ 3D locations. 0pt uses basic geometry calculations and requires 2 assumptions to hold. First, the object of interest must be on the ground. We can then assume that the lower horizontal border of the object’s bounding box is the point where the object touches the ground. Second, we can identify the equation of the plane that the object of interest is on. In Spatialyze, the plane $z=0$ is the ground, although our algorithm can be modified to use with any plane. The middle point of the lower border of each object’s 2D bounding box represents the point that it touches the ground (red star in Fig. 7).
@@ 6.3.2. Algorithm
0pt performs 2 steps to estimate 3D location of an object from its 2D bounding box. First, based on the second assumption, 0pt finds an equation that represents a Vector of Possible 3D Locations of an object. Second, based on the first and third assumptions, 0pt finds an Exact 3D Location from the vector that intersects with the ground.
Vector of Possible 3D Locations. Using Eq. 5, we convert any 2D point in pixel coordinate system to a 3D point in world coordinate system, if we know $z_{c}$. Since we do not know $z_{c}$ in this case, we can derive a parametric equation representing all possible 3D points (red line in Fig. 7) converted from the 2D point. Let $d$ be the unknown distance of the object from the camera.
@@
Exact 3D Location. From our assumptions, we know that the bottom line of the object’s bounding box touches the ground. Therefore, $(x,y,z)$ also touches the ground. Because the ground is $z=0$ in our case, we can solve the Eq. 7 for $d$. Once we know $d$, we can solve the parametric Eq. 7 for $x$ and $y$ (blue star in Fig. 7).
@@ 6.3.3. Benefits and Limitations
0pt relies on the assumptions mentioned, which do not always hold, especially, the first one. Nevertheless, Spatialyze can analyze users’ filter predicates and tell whether the types of objects of interest can be assumed to touch the ground and only applies the optimization then. For example, “car,” “bicycle,” or “human” touch the ground, while “traffic light” does not. In addition, 0pt is 192$\times$ faster compared than Monodepth2 on average. If one of the estimated 3D locations of an object in a video frame ends up being behind the camera, which means that the object does not touch the ground, Spatialyze can fall back to use Monodepth2 for that frame. Otherwise, 0pt provides 99.48% runtime reduction on average for object 3D location estimation and 52% for the video processing.
Figure 7. 0pt recovers the 3D location (blue star) of a car from its 2D location (red star).
@@ 6.4. Exit Frame Sampler
Object trajectories are computed in the object tracker step (§ 5.2.2) using an object tracking algorithm. Such algorithms, including that used in Spatialyze, perform tracking-by-detection (Leal-Taixé, 2014). An object detector identifies objects in video frames but doesn’t link detections across frames to represent the same object. The object tracker associates detections in consecutive frames, creating a chain of associated detections that forms an object trajectory, such as a Movable Object. However, not all detections are necessary in an object trajectory. For instance, a car travels in a straight line from frame 1 to 3, where it is at location (0, 1), (0, 2), (0, 3) at frame 1, 2, 3. The in-between detection (0, 2) at frame 2 is unnecessary as the detections at frame 1 and frame 3 produce the same trajectory.

### v4

@@ 6.1.2. Algorithm
Pixel to Camera Coordinate System. The equation for converting a 3D point $(x_{c},y_{c},z_{c})$ from camera coordinate system333 Camera Coordinate System is a 3-dimensional Euclidean space from the perspective of a camera. The origin point is at the camera position. Z-axis and X-axis point toward the front and right of the camera. The Y-axis points downward from the camera. to a 2D point $(x_{p},y_{p})$ in pixel coordinate system is:
@@
where $it_{i}$ represents the camera intrinsic, a set of parameters representing Camera’s focal lengths ($f_{x},f_{y}$), skew coefficient ($s$), and optical center in pixels ($x_{0},y_{0}$). The intrinsic is specific to each camera, independent of its location and rotation. It converts 3D points from a camera perspective to 2D points in pixels.
@@
(3) $\begin{split}\left[\begin{smallmatrix}x_{c}\\ y_{c}\\ z_{c}\\ 1\end{smallmatrix}\right]&=\left[\begin{smallmatrix}\begin{array}[]{c|c}\text{\small\emph{it}}_{\text{\scriptsize\em i}}^{\text{\scriptsize-1}}&\text{\small{0}}\\ \hline\cr\text{\small{0}}&\text{\small{1}}\end{array}\end{smallmatrix}\right]\times\left[\begin{smallmatrix}x_{p}z_{c}\\ y_{p}z_{c}\\ z_{c}\\ 1\end{smallmatrix}\right]\\[-8.0pt] &=C\times\left[\begin{smallmatrix}x_{p}z_{c}&y_{p}z_{c}&z_{c}&1\end{smallmatrix}\right]^{\top},\text{where }C=\left[\begin{smallmatrix}\begin{array}[]{c|c}\text{\small\emph{it}}_{\text{\scriptsize\em i}}^{\text{\scriptsize-1}}&\text{\small{0}}\\ \hline\cr\text{\small{0}}&\text{\small{1}}\end{array}\end{smallmatrix}\right]\end{split}$
Camera to World Coordinate System. For this conversion, we set up $R$, a $3\times 3$ matrix, representing the 3D rotation of the camera that is derived from the camera’s rotation quaternion. And, $t$ is a 3D vector, representing the camera’s translation. Then, a $3\times 4$ extrinsic matrix $[R|t]$ converts 3D points from camera coordinate system $(x_{c},y_{c},z_{c})$ to 3D points $(x,y,z)$ in world coordinate system:
Pixel to World Coordinate System. Combining the previous 2 conversions in Eq. 3 and Eq. 4 converts 2D points from a pixel coordinate system to 3D points in the world coordinate system:
4 corners of a video frame in the World Coordinate System. To convert the 4 corners of a video frame from pixel to world coordinates, we modify Eq. 5 where we substitute $z_{c}$ with $d$, and $x_{p}$ and $y_{p}$ with actual pixel values representing the 4 corners of the video frame; $w$ and $h$ are the width and height in pixels of the video frame, and each of $B,C,D,E$ is a $3\times 1$ vector representing its 3D location.
@@
Finally, the Road Visibility Pruner uses the users’ filter predicates to determine if there is any Geographic Construct of interest visible in the frame. Examining the contains predicates in the filter, it assigns each contains to $True$ if its first argument is in visibleGeogs and $False$ otherwise. If the value of the transformed filter predicates is $False$, then the Road Visibility Pruner prunes out that frame.
@@ 6.1.3. Benefits and Limitations
The Road Visibility Pruner has an insignificant overhead runtime compared to the runtime it can save. Our experiments show that the Road Visibility Pruner takes 0.1% of the video processing runtime to execute while saving up to 19.6% of the video processing runtime. So in the worst case that Road Visibility Pruner cannot prune out any video frames, the runtime increase is still negligible. However, having the Road Visibility Pruner in the video processing plan may cause tracking accuracy to drop in some scenarios. Using § 3 as an example, if a car starts at an intersection, exits the intersection, and enters another intersection. When the car is not in any intersection, the camera may not capture any frames containing an intersection. The Road Visibility Pruner will prune out the period of video frames where no intersection is visible to the camera. The object tracker might consider the car before leaving the intersection and the car after entering another intersection to be two different cars.
[…]
The illustrations of the side and top view of the camera’s viewable space (in pyramid shape), where points A represents the camera position, and points B, C, D, and E represent the top-left, top-right, bottom-right, bottom-left of the frame at d meters in front of the camera. The left figure shows the camera’s viewable area intersecting with a polygon representing a road intersection. The right figure shows the camera’s viewable area, which does not intersect with any polygon representing a road intersection.
[…]
@@ 6.2.1. High-Level Concepts
Spatialyze only needs to track objects of interest as stated in the workflow. The Object Type Pruner prunes out irrelevant objects from being tracked. We analyze users’ filter predicates to look for types of objects that are necessary for computing the workflow outputs. Using the predicates in § 3 as an example, the Object Type Pruner partially executes the line 3, where users only look for cars and trucks in the workflow outputs. The trajectories of other objects (e.g., humans) will not appear in the final workflow outputs anyway, so tracking them is unnecessary. Therefore, we only need to input objects detected as cars or trucks into the object tracker to reduce its workload.
@@ 6.2.2. Benefits and Limitations
The Object Type Pruner reduces the object tracker’s runtime. It has low overhead runtime of 0.06% and does not require geospatial metadata but can reduce up to 69% of the tracking runtime, which is 26% of the video processing time.
@@
In our video processing plan, Spatialyze estimates an object’s 3D location using its distance from the camera and calculated bounding box. Specifically, we use YOLOv5 (Jocher et al., 2022) to estimate the bounding box and Monodepth2 (Godard et al., 2019) to estimate the object’s distance from the camera in our baseline plan. Monodepth2 only requires monocular images to estimate each object’s distance from the camera; therefore, it is flexible in general use cases. Spatialyze can always use Monodepth2 to estimate any object’s distance from the camera in a video frame. However, this algorithm is slow and can take up to 48% portion of the total baseline video processing runtime. Therefore, we explored an alternative algorithm that leverages existing geospatial metadata to speed up 3D object location estimation.
@@ 6.3.1. High-Level Concepts
0pt leverages Camera to estimate an object’s 3D locations from its 2D bounding box. It is a substitution for Monodepth2 to estimate objects’ 3D locations. 0pt uses basic geometry calculations and requires 2 assumptions to hold. First, the object of interest must be on the ground. We can then assume that the lower horizontal border of the object’s bounding box is where the object touches the ground. Second, we can identify the equation of the plane that the object of interest is on. In Spatialyze, the plane $z=0$ is the ground, although our algorithm can be modified to be used with any plane. The middle point of the lower border of each object’s 2D bounding box represents the point that it touches the ground (red $\star$ in Fig. 3).
@@ 6.3.2. Algorithm
0pt performs 2 steps to estimate the 3D location of an object from its 2D bounding box. First, based on the second assumption, 0pt finds an equation that represents a Vector of Possible 3D Locations of an object. Second, based on the first and second assumptions, 0pt finds an Exact 3D Location from the vector that intersects with the ground.
Vector of Possible 3D Locations. Using Eq. 5, we convert any 2D point in the pixel coordinate system to a 3D point in the world coordinate system, if we know $z_{c}$. Since we do not know $z_{c}$ yet, we can derive a parametric equation representing all possible 3D points (red line in Fig. 3) converted from the 2D point. Let $d$ be the unknown distance of the object from the camera.
@@
Exact 3D Location. We assume that the bottom line of the object’s bounding box touches the ground; therefore, $(x,y,z)$ also touches the ground. Because the ground is $z=0$ in our case, we can solve the Eq. 7 for $d$. Once we know $d$, we can solve the parametric Eq. 7 for $x$ and $y$ (blue $\star$ in Fig. 3).
@@ 6.3.3. Benefits and Limitations
0pt relies on the assumptions mentioned, which do not always hold, especially, the first one. Nevertheless, Spatialyze can analyze users’ filter predicates and tell whether the types of objects of interest can be assumed to touch the ground, applying the optimization only then. For example, “car,” “bicycle,” or “human” touch the ground, while “traffic light” does not. In addition, 0pt is 192$\times$ faster compared than Monodepth2 on average. If one of the estimated 3D locations of an object in a video frame ends up being behind the camera, which means that the object does not touch the ground, Spatialyze can fall back to using Monodepth2 for that frame. Otherwise, 0pt provides 99.48% runtime reduction on average for object 3D location estimation and 52% for the video processing.
[…]
A 3D illustration of geometry-based 3D location estimation, showing relationships between the camera, the 2D car in the image frame, and the actual car location
Figure 3. From a car’s 2D location ($\star$) to its 3D location ($\star$).
@@ 6.4. Exit Frame Sampler
Common object tracking algorithms (Du et al., 2023; Wojke et al., 2017; Bewley et al., 2016), including that used in Spatialyze in the object tracker (§ 5.2.2), performs tracking-by-detection (Leal-Taixé, 2014). An object detector identifies objects in video frames but does not link detections across frames to represent the same object. The object tracker associates detections in consecutive frames, creating a chain of associated detections that forms an object trajectory, such as a Movable Object. However, not all detections are necessary in an object trajectory. For instance, a car travels in a straight line from frame 1 to 3, where it is at location (0, 1), (0, 2), (0, 3) at frame 1, 2, 3. The in-between detection (0, 2) is unnecessary as the detections at frame 1 and frame 3 produce the same trajectory.